OpenStudy (anonymous):

8x−9=3x=16 What is the value of x?

OpenStudy (xmissalycatx):

Hey again! :) I was wondering if the equation was actually supposed to say = 2 times or if that was a typo?

OpenStudy (anonymous):

oh, the second = is supposed to be a +

OpenStudy (anonymous):

First,combine the like terms

OpenStudy (anonymous):

\[8x-9=3x+16\]

OpenStudy (xmissalycatx):

Lol okay thank you.

OpenStudy (anonymous):

\[8x-3x=16+9\]

OpenStudy (anonymous):

\[5x=25\]Then,divide both sides by 5

OpenStudy (xmissalycatx):

so just combine your like terms! :) 8x - 3x would be 5x and 16 + 9 would be 25 So now you have 5x = 25 So just divide BOTH sides by 5

OpenStudy (anonymous):

\[\frac{ 5x }{ 5 }=\frac{ 25 }{ 5 }\]\[x=?\]solve it @Girlie310

OpenStudy (anonymous):

can i have some help?'

OpenStudy (anonymous):

sure

OpenStudy (xmissalycatx):

@Girlie310 still there? :P